OpenStudy (anonymous):

4x – 2(–2x – 3) = 10x A. –8 B. 3 C. no solution D. all real numbers

OpenStudy (anonymous):

\[4x – 2(–2x – 3) = 10x \]Distribute the -2. \[-2 \times -2x = 4x \quad and -2 \times -3 = 6\] \[4x + 4x + 6 = 10x \]Combine like terms.\[8x + 6 = 10x\]Subtract the 8x from both sides.\[8x -8x+ 6 = 10x - 8x\]\[6 = 2x\]Divde both sides by 2.\[x = 6 \div 2 \]

OpenStudy (anonymous):

what website is that

OpenStudy (anonymous):

Oh and, welcome to OpenStudy!

OpenStudy (anonymous):

What website is what?

OpenStudy (anonymous):

the where you just got the answer

OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2) A. 30 B. –15 C. no solution D. all real numbers

OpenStudy (anonymous):

I didn't get the answer from a website. I solved it myself. (: